# Global calibration of large-scale photonic integrated circuits

### v1

@@ III.2 Experimental results
Figure 5: Experimental calibration results and validation of the QW PIC. (a) and (b) depict the predicted values of the parameters $\bm{a}$ and $\bm{b}$ in the phase-current relation for each phase shifter, respectively. (c) Predicted values of the reflectivity $\sin^{2}\bm{\theta}$ for individual on-chip beamsplitter. (d) Predicted relative coupling efficiency $\bm{\eta}$ between the output ports. (e) Histogram of the measured values of the L1-norm distances $d(\bm{\lambda}^{\ast})$, which represent the difference between the test dataset’s output light probability distributions $\{\bm{P}^{\text{test}}_{\text{exp}}\}_{n}$ and the fully trained network’s predictions $\{\bm{P}_{\text{ML}}(\bm{\lambda}^{\ast})\}_{n}$. (f) The measured probability distribution (blue bars) for 12-step Hadamard QWs. The red bars mean the theoretical expectation’s distribution, taking into account the unbalanced reflectivities of the waveguide beamsplitters shown in (c).
@@
Herein, the calibration of the QW PIC is implemented and tested by training the custom network directly on actual experimental datasets $\{\bm{P}_{\text{exp}}\}_{n}$. In order to generate the experimental training and test datasets, we randomly generate 1500 sets of current distributions $\{\bm{I}\}_{n}$ ranging from 0 mA to 7 mA imposed on the phase shifters on our silicon PIC. We obtain 1500 sets of experimental probability distributions $\{\bm{P}_{\text{exp}}\}_{n}$ after normalizing the measured light intensity over all output ports. The experimental datasets are divided into two parts according to an 85:15 ratio, where 1275 sets of datasets are used to train the custom network, and the remaining 225 sets of datasets are used to test the trained network. With these experimental training sets, the parameters $\bm{\lambda}=\{\bm{a},\bm{b},\bm{\theta},\bm{\eta}\}$ of the custom network in Fig. 1 are then pre-trained.
@@
The fully trained network can characterize the phase-current relation, assumed to be of the form $\phi=aI^{2}+b$, the reflectivity parameter $\bm{\theta}$ of beamsplitters, and the relative output coupling efficiency $\bm{\eta}$ between the output ports. The characterization results for 65 phase shifters, 78 beamsplitters and 20 output ports are shown in Figs. 5(a)-5(d). Note that we only experimentally detect 20 output ports in the middle, neglecting two output ports at the top and bottom positions, as the output light intensities at these four ports are not affected by the phase shifter settings. The phase shifter in the upper right corner of the QW PIC is also omitted because it has no impact on the output light intensity. Due to the influence of the chip fabrication errors, the coefficients $\bm{a}$ and $\bm{b}$ in Figs. 5(a) and (b) for each phase shifter are often not the same, and the reflectivity of some beamsplitters in Fig. 5(c) deviates from the ideal value of 0.5.
@@
After completing the calibration of the photonic chip using the training datasets, we utilize the experimental test datasets to further verify the accuracy of the calibration. We apply the remaining 225 sets of distributions of random currents, and compare the experimentally measured distributions $\{\bm{P}^{\text{test}}_{\text{exp}}\}_{n}$ with the predicted probability distributions $\{\bm{P}_{\text{ML}}(\bm{\lambda}^{\ast})\}_{n}$ of the trained network with the optimal hyperparameters $\bm{\lambda}^{\ast}$. The difference between the two distributions for each test sample $n$ is defined by L1-norm distance: $d(\bm{\lambda}^{\ast})=\frac{1}{2}|\bm{P{{}_{\text{ML}}}}(\bm{\lambda}^{\ast})-\bm{P}^{\text{test}}_{\text{exp}}|$. We observe that the distance values for $77\%$ of the test samples are less than 0.05, as shown in Fig. 5(e). The average distance for all test samples is $\bar{d}=0.037\pm 0.016$, indicating that our calibration results for the phase shifters and beamsplitters are in good agreement with the actual physical process. The slight deviation suggests the presence of additional chip imperfections that are not considered by our custom network, such as the non-dominant thermal crosstalk effect [54]. The error bar of $\bar{d}$ represents one standard deviation.
@@
After verifying the performance on numerous random validation samples using the calibrated photonic chip, we proceed to implement the conventional 1D Hadamard QWs, where the coin operators are fixed as the Hadamard gate and independent of time and position. We inject 1550 nm light into the lower path of the first waveguide beamsplitter followed by setting the phase of the first phase shifter to $\pi/2$. Therefore, we prepare a balanced superposition initial state between two paths with a phase difference of $\pi/2$ through one-step QW dynamics [55]. The phases of the remaining phase shifters are set to 0. We measure the light intensity after the 12-step QWs, and the experimental probability distribution is depicted in Fig. 5(f). The imbalance between the two major peak probabilities at the third and eighteenth output ports is mainly derived from the difference in the reflectivities of different waveguide beamsplitters shown in Fig. 5(c). To confirm this, we numerically calculate the theoretical probability distribution of the 12-step QWs model considering the calibrated reflectivities and then compare it with the experimentally observed result. The resulting distance between the two distributions is $d=0.046$, indicating their good agreement. The unbalanced distribution induced by the static fabrication error in beamsplitters could be compensated by adjusting the phase shift of phase shifters to achieve high-fidelity unitary dynamics [56].
@@ IV Conclusion
In conclusion, we introduce a white-box machine learning method for the global calibration of integrated photonic circuits. To achieve it, we construct a network with custom layers to model the physical process of the chip and train it using the gradient-descent-based tool, Adam. We experimentally validate our calibration method using a photonic QW chip with a circuit depth of 12. The trained network can globally calibrate the nonlinear phase-current relations of the thermo-optical phase shifters and the reflectivities of beamsplitters on the QW chip, which efficiently bypasses the error accumulation problem associated with the traditional one-by-one calibration method. Consequently, a low average distance of 0.037 between the measured chip distributions and the network predictions is achieved.
Our calibration method does not require a meticulous design of the decoupling scheme to calibrate phase shifters one by one, making it suitable for photonic chips with a prior physical process. Moreover, this method offers high degrees of freedom, allowing any parameters involved in the physical process to be added to the custom network and learned through gradient-descent-based optimization. For instance, thermal crosstalk parameters between phase shifters could be further characterized by computing the nondiagonal terms in a given matrix phase-current relation. Thus, our method shows promise as a potent tool with potential applications in large-scale chip calibration and optimal control problems [57, 58, 59, 60].
@@ Acknowledgements.
J.-H. Zheng and Q.-Q. Wang contributed equally to this paper. This work was supported by Innovation Program for Quantum Science and Technology (Nos. 2021ZD0301200, 2021ZD0301500, 2021ZD0303200), National Natural Science Foundation of China (Nos. 12022401, 62075207, 62275240, 11874343, 12104433, 11821404, 12204468), China Postdoctoral Science Foundation (No. 2021M703108), Fundamental Research Funds for the Central Universities (Nos. WK2470000030, WK2030000081), CAS Youth Innovation Promotion Association (No. 2020447), and Research and Development Program of Anhui Province (No. 2022b1302007). We thank Hefei Guizhen Chip Technologies Co., Ltd. for collaborating to develop the multi-channel current/voltage source.

### v2

@@ III.1 Experimental setup
Figure 6: Experimental calibration results. (a) and (b) depict the predicted values of the parameters $\bm{a}$ and $\bm{b}$ in the phase-current relation for each phase shifter, respectively. (c) Predicted values of the reflectivity $\sin^{2}\bm{\theta}$ for individual on-chip beamsplitter. (d) Predicted relative transfer efficiency $\bm{\alpha}$ between the beamsplitters. (e) Predicted relative coupling efficiency $\bm{\eta}$ between the output ports. (f) Histogram of the measured values of the L1-norm distance $d(\bm{\lambda}^{\ast})$, which represent the difference between the test dataset’s output light probability distributions $\{\bm{P}^{\text{test}}_{\text{exp}}\}_{n}$ and the trained network’s predictions $\{\bm{P}_{\text{ML}}(\bm{\lambda}^{\ast})\}_{n}$.
@@ III.2 Experimental results
Herein, the calibration of the QW PIC is implemented and tested by training the custom network directly on actual experimental datasets $\{\bm{P}_{\text{exp}}\}_{n}$. In order to generate the experimental training and test datasets, we randomly generate 1500 sets of current distributions $\{\bm{I}\}_{n}$ ranging from 0 mA to 7 mA imposed on the phase shifters on our silicon PIC. We obtain 1500 sets of experimental probability distributions $\{\bm{P}_{\text{exp}}\}_{n}$ after normalizing the measured light intensity over all output ports. The experimental datasets are divided into two parts according to an 85:15 ratio, where 1275 sets of datasets are used to train the custom network, and the remaining 225 sets of datasets are used to test the trained network. With these experimental training sets, the parameters $\bm{\lambda}=\{\bm{a},\bm{b},\bm{\theta},\bm{\alpha},\bm{\eta}\}$ of the custom network in Fig. 1 are then pre-trained.
The trained network can characterize the phase-current relation, assumed to be of the form $\phi=aI^{2}+b$, the reflectivity parameter $\bm{\theta}$ and the relative transfer efficiency $\bm{\alpha}$ of BSs, as well as the relative output coupling efficiency $\bm{\eta}$ between the output ports. Here, we provide the relative transfer efficiency between all BSs and the relative coupling efficiency between all output ports because we utilize the normalized probability distributions to calibrate the QW PIC. Even so, we only need to measure the absolute transfer efficiency for one BS and the absolute coupling efficiency for one output port. The efficiencies of all the remaining BSs and output ports can then be directly calculated.
@@
The characterization results for 65 phase shifters, 78 beamsplitters and 20 output ports are shown in Figs. 6(a)-6(e). Note that we only experimentally detect 20 output ports in the middle, neglecting two output ports at the top and bottom positions, as the output light intensities at these four ports are not affected by the phase shifter settings. The phase shifter in the upper right corner of the QW PIC is also omitted because it has no impact on the output light intensity. Due to the influence of the chip fabrication errors, the coefficients $\bm{a}$ and $\bm{b}$ in Figs. 6(a) and 6(b) for each phase shifter are often not the same, and the reflectivity of some beamsplitters in Fig. 6(c) deviates from the ideal value of 0.5.
Figure 7: Validation of the QW PIC. (a) The measured probability distribution $\bm{P}^{\text{QW}}_{\text{exp}}$ (blue bars) for 12-step Hadamard QWs. The red bars mean the distribution $\bm{P}_{\text{ML}}(\bm{\lambda}^{\ast})$ predicted by the trained network. (b) and (c) display histograms of the measured values of infidelity and purity error for 100 output quantum states after 9-step QW dynamics, respectively.
@@
After completing the calibration of the photonic chip using the training datasets, we utilize the experimental test datasets to further verify the accuracy of the calibration. We apply the remaining 225 sets of distributions of random currents, and compare the experimentally measured distributions $\{\bm{P}^{\text{test}}_{\text{exp}}\}_{n}$ with the predicted probability distributions $\{\bm{P}_{\text{ML}}(\bm{\lambda}^{\ast})\}_{n}$ of the trained network with the optimal hyperparameters $\bm{\lambda}^{\ast}$. The difference between the two distributions for each test sample $n$ is defined by L1-norm distance: $d(\bm{\lambda}^{\ast})=\frac{1}{2}|\bm{P{{}_{\text{ML}}}}(\bm{\lambda}^{\ast})-\bm{P}^{\text{test}}_{\text{exp}}|$. We observe that the distance values for $85\%$ of the test samples are less than 0.05, as shown in Fig. 6(f). The average distance for all test samples is $\bar{d}=0.033\pm 0.018$, indicating that our calibration results for the phase shifters and beamsplitters are in good agreement with the actual physical process. The slight deviation suggests the presence of additional chip imperfections that are not considered by our custom network, such as the non-dominant thermal crosstalk effect [54]. The error bar of $\bar{d}$ represents one standard deviation.
@@
After verifying the performance on numerous random validation samples using the calibrated photonic chip, we proceed to implement the conventional 1D Hadamard QWs, where the coin operators are fixed as the Hadamard gate and independent of time and position. We inject 1550 nm light into the lower path of the first waveguide beamsplitter followed by setting the phase of the first phase shifter to $\pi/2$. Therefore, we prepare a balanced superposition initial state between two paths with a phase difference of $\pi/2$ through one-step QW dynamics [55]. The phases of the remaining phase shifters are set to 0. We measure the light intensity coupled to the optical fiber arrays, and the experimental probability distribution $\bm{P}^{\text{QW}}_{\text{exp}}$ for a 12-step QW is obtained using the calibrated fiber-coupling efficiency for compensation, as depicted in Fig. 7(a). The resulting L1-norm distance with respect to the QW distribution predicted by the trained network is $d=0.057$, indicating their good agreement.
Beyond the position distributions, we then perform the quantum state tomography for the calibrated QW PIC with inherent high-dimensional structures [56, 57, 58, 59]. To achieve this, in addition to the computational basis |$\ket{v}$, extra measurement bases need to be introduced to fully characterize the output quantum state. We divide the QW PIC into two sections: the first part, with a circuit depth of 9, is used to implement arbitrary 9-step QW dynamics, while the second part, with a circuit depth of 3, serves as a high-dimensional measurement device [60, 61, 62, 63]. For each 9-step QW dynamics, we impose 14 sets of current distributions into the measurement device and measure the output probability distributions. Here, we perform standard maximum-likelihood reconstruction with the estimator taking the form ${\rho_{\text{m}}=TT^{\dagger}/\text{Tr}(TT^{\dagger})}$, where $T$ is a lower triangular matrix [64]. The quantum states after the 9-step QW can be reconstructed by maximizing the likelihood of 14 sets of experimental distributions. We arbitrarily generate 100 types of QW dynamics, and the histograms of the infidelity and purity error of the reconstructed quantum states are displayed in Figs. 7(b) and 7(c), respectively. The corresponding average values of the infidelity and purity error are $0.025\pm 0.012$ and $0.042\pm 0.037$, respectively, indicating that the experimental states align well with the theoretical expectations.
@@ IV Conclusion
In conclusion, we introduce a global calibration method for the integrated photonic circuits. To achieve this, we construct a network with custom layers to model the physical process of the chip and train it using the gradient-descent-based tool, Adam. We experimentally validate our calibration method using a photonic QW chip with a circuit depth of 12. The trained network can globally calibrate the nonlinear phase-current relations of the thermo-optical phase shifters and the reflectivities of beamsplitters on the QW chip, which efficiently bypasses the error accumulation problem associated with the traditional one-by-one calibration method. Consequently, a low average distance of 0.033 between the measured chip distributions and the network predictions is achieved.
Our calibration method does not require a meticulous design of the decoupling scheme to calibrate phase shifters one by one, making it suitable for photonic chips with a prior physical structure. Moreover, this method offers high degrees of freedom, allowing any parameters involved in the physical process to be added to the custom network and learned through gradient-descent-based optimization. For instance, thermal crosstalk parameters between phase shifters could be further characterized by computing the nondiagonal terms in a given matrix phase-current relation. Thus, our method shows promise as a potent tool with potential applications in large-scale chip calibration and optimal control problems [65, 66, 67, 68].
@@ Acknowledgements.
J.-H. Zheng and Q.-Q. Wang contributed equally to this paper. This work was supported by Innovation Program for Quantum Science and Technology (Nos. 2021ZD0301200, 2021ZD0301500, 2021ZD0303200), National Natural Science Foundation of China (Nos. 12022401, 62075207, 62275240, 11874343, 12104433, 11821404, 12204468), China Postdoctoral Science Foundation (No. 2024M753083), National Postdoctoral Program for Innovative Talents (No. BX20240353), Students’ Innovation and Entrepreneurship Foundation of USTC (No. SC5290005208), Fundamental Research Funds for the Central Universities (Nos. WK2470000030, WK2030000081), CAS Youth Innovation Promotion Association (No. 2020447), and Research and Development Program of Anhui Province (No. 2022b1302007). We thank Hefei Guizhen Chip Technologies Co., Ltd. for collaborating to develop the multi-channel current/voltage source.